OpenStudy (anonymous):

Solve equation by completing the square. 1) p^2+14p+9=-5

OpenStudy (lgbasallote):

transpose 9 first p^2 + 14p + ___ = -14 to get the third term of the left hand side, divide the numerical coefficient of the second term by two then square the quotient so the third term is (14/2)^2 = 7^2 = 49 so p^2 + 14p + 49 = -14 + 49 (p+7)^2 = 35 square root... p + 7 = +/- sqrt 35 p = 7 +/- sqrt 35

OpenStudy (anonymous):

so that would be it?

OpenStudy (anonymous):

is there any way that it would be faster? or simple ?